What is the percent by mass of a solution that contains 25.00 g of NaCl in 150.0 g water?




1
Expert's answer
2021-02-24T03:37:43-0500

mass% =100%* m(NaCl)/ (mNaCl) + m(H2O))= 25/175=14.29%


Answer: 14.29%